# A modern perspective on rational homotopy theory

### v1

@@ 7. The family of model structures on $1$-reduced simplicial sets
Let $M$ be a multiplicative subset of $\mathbb{Z}$. We left Bousfield localize the model category ${\mathcal{SS}\mathrm{ets}}_{1}$ to get a model category $\mathcal{L}_{M^{-1}\mathbb{Z}}{\mathcal{SS}\mathrm{ets}}_{1}$ whose weak equivalences are the $M$-local homotopy equivalences. In particular, we obtain a model category $\mathcal{L}_{\mathbb{Q}}{\mathcal{SS}\mathrm{ets}}_{1}$ whose weak equivalences are the rational homotopy equivalences. Thus, we recover the family of model structures on $1$-reduced simplicial sets from [Qui69, Part II, Theorem 2.2]. These model structures inherit the generating cofibrations and the simplicial structure that we found for ${\mathcal{SS}\mathrm{ets}}_{1}$.
@@ 7.1. Left Bousfield localization
We give an overview of left Bousfield localization; a reference is [Hir03, Chapters 1-4]. Let $\mathcal{M}$ be a simplicial model category and $T$ be a set of morphisms in $\mathcal{M}$.
@@
    An object $K$ of $\mathcal{M}$ is $T$-local if $K$ is fibrant in $\mathcal{M}$ and, for every morphism $\varphi\colon A\rightarrow B$ in $T$, the map
    $\varphi^{\ast}\colon\operatorname{Map}_{\mathcal{M}}\left(B,K\right)\rightarrow\operatorname{Map}_{\mathcal{M}}\left(A,K\right)$
    is a weak homotopy equivalence.
@@
    A morphism $f\colon X\rightarrow Y$ in $\mathcal{M}$ is a $T$-local equivalence if, for every $T$-local object $K$, the map
    $f^{\ast}\colon\operatorname{Map}_{\mathcal{M}}\left(Y,K\right)\rightarrow\operatorname{Map}_{\mathcal{M}}\left(X,K\right)$
    is a weak homotopy equivalence.
@@
    A $T$-localization of an object $X$ of $\mathcal{M}$ is a $T$-local equivalence $f\colon X\rightarrow X_{T}$ with $T$-local target.
Example 7.2.
Every morphism in $T$ and every weak equivalence in $\mathcal{M}$ is a $T$-local equivalence.
Theorem 7.3 ([Bar10, Theorem 4.7]).
Let $\mathcal{M}$ be a combinatorial, left proper, and simplicial model category. Let $T$ be a set of morphisms in $\mathcal{M}$. There is a model structure $\mathcal{L}_{T}\mathcal{M}$ on the underlying category of $\mathcal{M}$, called the left Bousfield localization of $\mathcal{M}$ at $T$, in which a morphism $f\colon X\rightarrow Y$ is
@@
The left Bousfield localization of $\mathcal{M}$ at $T$ enlarges, in a systematic fashion, the class of weak equivalences of $\mathcal{M}$ to encompass the morphisms in the set $T$. Observe that $\mathcal{L}_{T}\mathcal{M}$ inherits the generating cofibrations of $\mathcal{M}$, but it generally does not also inherit the generating acyclic cofibrations of $\mathcal{M}$ as, in general, the class of acyclic cofibrations in $\mathcal{M}$ is properly contained in the class of acyclic cofibrations in $\mathcal{L}_{T}\mathcal{M}$.
@@ 7.2. The set for the left Bousfield localization
$\left\{S^{n}\rightarrow S^{n}_{M^{-1}\mathbb{Z}}\mid n\geq 2\right\}$
in order to left Bousfield localize the model category ${\mathcal{SS}\mathrm{ets}}_{1}$ at that set.
A sensible first suggestion is to start with the simplicial $n$-sphere $S_{\mathrm{simp}}^{n}$ and use it to emulate the construction of $S^{n}_{M^{-1}\mathbb{Z}}$ in the $1$-reduced setting. This approach runs into a problem: the construction of $S^{n}_{M^{-1}\mathbb{Z}}$ involves self-maps of $S^{n}$ of arbitrary degree, whereas the simplicial $n$-sphere $S_{\mathrm{simp}}^{n}$ only has the trivial self-maps.
@@
Let $f\colon S_{\mathrm{simp}}^{n}\rightarrow S_{\mathrm{simp}}^{n}$ be a self-map of $S_{\mathrm{simp}}^{n}$. Let $x$ denote the unique 00-simplex of $S_{\mathrm{simp}}^{n}$ and let $\epsilon$ denote the non-degenerate $n$-simplex of $S_{\mathrm{simp}}^{n}$. There are two possibilities for the $n$-simplex $f_{n}(\epsilon)$ of $S_{\mathrm{simp}}^{n}$.
  1. (1)
    If $f_{n}(\epsilon)=\epsilon$, then $f$ is the identity map of $S_{\mathrm{simp}}^{n}$.
  2. (2)
    If $f_{n}(\epsilon)$ is the iterated degeneracy of $x$, then $f$ is the constant self-map of $S_{\mathrm{simp}}^{n}$.
Thus, we have to settle for an indirect route. We start with the CW inclusion
$S^{n}\rightarrow S^{n}_{M^{-1}\mathbb{Z}}.$
We apply the singular complex functor to obtain the inclusion of simplicial sets
Sing($\operatorname{Sing}(S^{n})\rightarrow\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}}).$
Then, we apply the $1$st Eilenberg subcomplex functor to obtain the inclusion of $1$-reduced simplicial sets
@@
To confirm that this map is the analog we are looking for, we study its geometric realization
$|E_{1}\operatorname{Sing}(S^{n})|\rightarrow|E_{1}\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}})|.$
In particular, we study the commutative diagram of CW complexes
${{|E_{1}\operatorname{Sing}(S^{n})|}}$${{|\operatorname{Sing}(S^{n})|}}$${{S^{n}}}$${{|E_{1}\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}})|}}$|Sing(${{|\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}})|}}$${{S^{n}_{M^{-1}\mathbb{Z}},}}$
where the horizontal maps in the left-hand square are CW subcomplex inclusions and the horizontal maps in the right-hand square are natural maps from the geometric realization-singular complex adjunction. We show that all the horizontal maps are weak homotopy equivalences.
@@
${{|E_{1}\operatorname{Sing}(S^{n})|}}$${{|\operatorname{Sing}(S^{n})|}}$${{S^{n}}}$${{|E_{1}\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}})|}}$|Sing(${{|\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}})|}}$${{S^{n}_{M^{-1}\mathbb{Z}},}}$
@@
Proposition 7.6 translates to the success of our approach. Thus, for a multiplicative subset $M$ of $\mathbb{Z}$, we left Bousfield localize ${\mathcal{SS}\mathrm{ets}}_{1}$ at the set of maps
$T_{M^{-1}\mathbb{Z}}=\left\{E_{1}\operatorname{Sing}(S^{n})\rightarrow E_{1}\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}})\mid n\geq 2\right\}.$
7.3. The family of model structures on $1$-reduced simplicial sets
@@
$T_{M^{-1}\mathbb{Z}}=\left\{E_{1}\operatorname{Sing}(S^{n})\rightarrow E_{1}\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}})\mid n\geq 2\right\}.$
In particular, we describe the $T_{M^{-1}\mathbb{Z}}$-local objects and the $T_{M^{-1}\mathbb{Z}}$-local equivalences in ${\mathcal{SS}\mathrm{ets}}_{1}$. To state our characterizations, we translate the definitions from Section 6.1 from the language of $1$-connected topological spaces to that of $1$-reduced simplicial sets using geometric realization.
@@
    A $1$-reduced simplicial set $K$ is $M$-local if $K$ is a Kan complex and $|K|$ is an $M$-local space.
@@
    An $M$-localization of a $1$-reduced simplicial set $X$ is an $M$-local homotopy equivalence $f\colon X\rightarrow X_{M^{-1}\mathbb{Z}}$ with $M$-local target.
@@
Example 7.8.
By Proposition 7.6, every map in $T_{M^{-1}\mathbb{Z}}$ is an $M$-localization. Note that Corollary 3.19 says that the target of every map in $T_{M^{-1}\mathbb{Z}}$ is a Kan complex.
The following proposition helps us transfer the rational homotopy theory of $1$-connected spaces from Section 6 to $1$-reduced simplicial sets.
Proposition 7.9 ([Hir03, Proposition 1.1.11]).
If $A$ and $X$ are pointed simplicial sets and $X$ is a Kan complex, then there is a weak homotopy equivalence
$\operatorname{Map}_{\mathcal{SS}\mathrm{ets}_{\ast}}\left(A,X\right)\rightarrow\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}\left(|A|,|X|\right)$
that is natural in $A$ and in $X$.
By a series of characterizations, we show that a $1$-reduced simplicial set is $T_{M^{-1}\mathbb{Z}}$-local if and only if it is $M$-local.
@@
Let $M$ be a multiplicative subset of $\mathbb{Z}$. Let $X$ be a $1$-reduced simplicial set. The following statements are equivalent.
@@
    The $1$-reduced simplicial set $X$ is $T_{M^{-1}\mathbb{Z}}$-local.
@@
    The $1$-reduced simplicial set $X$ is a Kan complex such that, for every $n\geq 2$, the map
    $\operatorname{Map}_{\mathcal{SS}\mathrm{ets}_{\ast}}\left(E_{1}\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}}),X\right)\rightarrow\operatorname{Map}_{\mathcal{SS}\mathrm{ets}_{\ast}}\left(E_{1}\operatorname{Sing}(S^{n}),X\right)$
    is a weak homotopy equivalence.
@@
    The $1$-reduced simplicial set $X$ is a Kan complex such that, for every $n\geq 2$, the map
    $\operatorname{Map}_{\mathcal{SS}\mathrm{ets}_{\ast}}\left(\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}}),X\right)\rightarrow\operatorname{Map}_{\mathcal{SS}\mathrm{ets}_{\ast}}\left(\operatorname{Sing}(S^{n}),X\right)$
    is a weak homotopy equivalence.
@@
    The $1$-reduced simplicial set $X$ is a Kan complex such that, for every $n\geq 2$, the map
    $\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}\left(|\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}})|,|X|\right)\rightarrow\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}\left(|\operatorname{Sing}(S^{n})|,|X|\right)$
    is a weak homotopy equivalence.
@@
    The $1$-reduced simplicial set $X$ is a Kan complex such that, for every $n\geq 2$, the map
    $\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}(S^{n}_{M^{-1}\mathbb{Z}},|X|)\rightarrow\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}(S^{n},|X|)$
    is a weak homotopy equivalence.
  6. (6)
    The $1$-reduced simplicial set $X$ is a Kan complex such that $|X|$ is an $M$-local space.
  7. (7)
    The $1$-reduced simplicial set $X$ is $M$-local.
Hence, $X$ is $T_{M^{-1}\mathbb{Z}}$-local if and only if $X$ is $M$-local.
@@
Firstly, (1) and (2) are equivalent by the definition of a $T_{M^{-1}\mathbb{Z}}$-local object in ${\mathcal{SS}\mathrm{ets}}_{1}$ combined with the characterization of fibrant objects in ${\mathcal{SS}\mathrm{ets}}_{1}$ from Theorem 3.17. Then, (2) and (3) are equivalent by the commutative diagram
${{\operatorname{Map}_{\mathcal{SS}\mathrm{ets}_{\ast}}\left(\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}}),X\right)}}$${{\operatorname{Map}_{\mathcal{SS}\mathrm{ets}_{\ast}}\left(E_{1}\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}}),X\right)}}$${{\operatorname{Map}_{\mathcal{SS}\mathrm{ets}_{\ast}}\left(\operatorname{Sing}(S^{n}),X\right)}}$${{\operatorname{Map}_{\mathcal{SS}\mathrm{ets}_{\ast}}\left(E_{1}\operatorname{Sing}(S^{n}),X\right),}}$
whose horizontal maps are weak homotopy equivalences by Proposition 7.6. Moreover, (3) and (4) are equivalent by the commutative diagram
${{\operatorname{Map}_{\mathcal{SS}\mathrm{ets}_{\ast}}\left(\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}}),X\right)}}$${{\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}\left(|\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}})|,|X|\right)}}$${{\operatorname{Map}_{\mathcal{SS}\mathrm{ets}_{\ast}}\left(\operatorname{Sing}(S^{n}),X\right)}}$${{\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}\left(|\operatorname{Sing}(S^{n})|,|X|\right),}}$
whose horizontal maps are weak homotopy equivalences by Proposition 7.9. Furthermore, (4) and (5) are equivalent by the commutative diagram
${{\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}(S^{n}_{M^{-1}\mathbb{Z}},|X|)}}$${{\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}\left(|\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}})|,|X|\right)}}$${{\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}(S^{n},|X|)}}$${{\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}\left(|\operatorname{Sing}(S^{n})|,|X|\right),}}$
whose horizontal maps are weak homotopy equivalences by Proposition 7.6. Also, (5) and (6) are equivalent by Proposition 6.9. Lastly, (6) and (7) are equivalent by the definition of an $M$-local simplicial set. ∎
By a similar series of characterizations, we establish that a map of $1$-reduced simplicial sets is a $T_{M^{-1}\mathbb{Z}}$-local equivalence if and only if it is an $M$-local homotopy equivalence. To that end, we need two auxiliary lemmas.
Lemma 7.11.
Let $M$ be a multiplicative subset of $\mathbb{Z}$. If $Z$ is an $M$-local pointed space, then $E_{1}\operatorname{Sing}(Z)$ is an $M$-local simplicial set.
Proof.
By Corollary 3.19, $E_{1}\operatorname{Sing}(Z)$ is a Kan complex. Then, Corollary 4.5 and Theorem 4.1(2) provide the weak homotopy equivalences
$|E_{1}\operatorname{Sing}(Z)|\rightarrow|\operatorname{Sing}(Z)|\rightarrow Z,$
so $|E_{1}\operatorname{Sing}(Z)|$ is $M$-local. Overall, $E_{1}\operatorname{Sing}(Z)$ is $M$-local, as required. ∎
@@
    For every $M$-local simplicial set $K$, the map
    $|f|^{\ast}\colon\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}(|Y|,|K|)\rightarrow\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}(|X|,|K|)$
    is a weak homotopy equivalence.
@@
    For every $M$-local pointed space $Z$, the map
    $|f|^{\ast}\colon\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}(|Y|,Z)\rightarrow\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}(|X|,Z)$
    is a weak homotopy equivalence.
@@
Suppose (1) holds. Let $Z$ be an $M$-local pointed space. Then, by Lemma 7.11, $E_{1}\operatorname{Sing}(Z)$ is an $M$-local simplicial set. There is a commutative diagram
${{\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}\left(|Y|,|E_{1}\operatorname{Sing}(Z)|\right)}}$${{\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}\left(|Y|,|\operatorname{Sing}(Z)|\right)}}$${{\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}(|Y|,Z)}}$${{\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}\left(|X|,|E_{1}\operatorname{Sing}(Z)|\right)}}$${{\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}\left(|X|,|\operatorname{Sing}(Z)|\right)}}$${{\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}(|X|,Z)}}$
whose horizontal maps are weak homotopy equivalences by Corollary 4.5 and Theorem 4.1(2) and whose left-hand vertical map is a weak homotopy equivalence by (1). Thus, the right-hand vertical map is a weak homotopy equivalence, which proves (2).
Conversely, suppose (2) holds. Let $K$ be an $M$-local simplicial set. Then, $|K|$ is an $M$-local pointed space, so
$|f|^{\ast}\colon\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}\left(|Y|,|K|\right)\rightarrow\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}\left(|X|,|K|\right)$
is a weak homotopy equivalence, which proves (1). ∎
We are now ready for the series of characterizations that describes the $T_{M^{-1}\mathbb{Z}}$-local equivalences.
@@
    The map $f$ is a $T_{M^{-1}\mathbb{Z}}$-local equivalence.
@@
    For every $T_{M^{-1}\mathbb{Z}}$-local simplicial set $K$, the map
    $f^{\ast}\colon\operatorname{Map}_{\mathcal{SS}\mathrm{ets}_{\ast}}(Y,K)\rightarrow\operatorname{Map}_{\mathcal{SS}\mathrm{ets}_{\ast}}(X,K)$
    is a weak homotopy equivalence.
@@
    For every $M$-local simplicial set $K$, the map
    $f^{\ast}\colon\operatorname{Map}_{\mathcal{SS}\mathrm{ets}_{\ast}}(Y,K)\rightarrow\operatorname{Map}_{\mathcal{SS}\mathrm{ets}_{\ast}}(X,K)$
    is a weak homotopy equivalence.

### v2

@@ 7. The family of model structures on $1$-reduced simplicial sets
Let $M$ be a multiplicative subset of $\mathbb{Z}$. We left Bousfield localize the model category ${\mathcal{SS}\mathrm{ets}}_{1}$ to get a model category $\mathcal{L}_{M^{-1}\mathbb{Z}}{\mathcal{SS}\mathrm{ets}}_{1}$ in which the weak equivalences are the $M$-local homotopy equivalences and the fibrant objects are the $M$-local Kan complexes. In particular, we obtain a model category $\mathcal{L}_{\mathbb{Q}}{\mathcal{SS}\mathrm{ets}}_{1}$ whose weak equivalences are the rational homotopy equivalences and whose fibrant objects are the rational Kan complexes. Thus, we recover the family of model structures on $1$-reduced simplicial sets from [33, Part II, Theorem 2.2] together with the characterization of fibrant objects in [33, Part II, Corollary 2.6]. These model structures inherit the generating cofibrations and the simplicial structure that we found for ${\mathcal{SS}\mathrm{ets}}_{1}$.
@@ 7.1. Left Bousfield localization
We first provide an overview of left Bousfield localization; a reference is [26, Chapters 1-4].
Let $\mathcal{M}$ be a simplicial model category; in particular, for two objects $X$ and $Y$, we have a function complex $\operatorname{Map}_{\mathcal{M}}(X,Y)$. Though one can left Bousfield localize model categories that are not simplicial using homotopy function complexes [26, Chapter 17], we only localize simplicial model categories in this paper, and our arguments are contingent upon our explicit understanding of the function complexes in our model categories.
Let $T$ be a set of morphisms in $\mathcal{M}$; in practice, it is important that $T$ is not a proper class.
@@
    An object $K$ of $\mathcal{M}$ is $T$-local if $K$ is fibrant in $\mathcal{M}$ and, for every morphism $\lambda\colon A\rightarrow B$ in $T$, the map $\operatorname{Map}_{\mathcal{M}}(\lambda,K)\colon\operatorname{Map}_{\mathcal{M}}(B,K)\rightarrow\operatorname{Map}_{\mathcal{M}}(A,K)$ is a weak equivalence.
@@
    A morphism $f\colon X\rightarrow Y$ in $\mathcal{M}$ is a $T$-local equivalence if, for every $T$-local object $K$, the map $\operatorname{Map}_{\mathcal{M}}(f,K)\colon\operatorname{Map}_{\mathcal{M}}(Y,K)\rightarrow\operatorname{Map}_{\mathcal{M}}(X,K)$ is a weak equivalence.
@@
    A $T$-localization of an object $X$ of $\mathcal{M}$ is a $T$-local equivalence with $T$-local target $f\colon X\rightarrow X_{T}$.
Remark 7.2.
Every morphism in $T$ and every weak equivalence in $\mathcal{M}$ is a $T$-local equivalence. Conversely, every $T$-local equivalence of $T$-local objects is a weak equivalence in $\mathcal{M}$ [26, Theorem 3.2.13(1)].
In light of our remark, the left Bousfield localization of $\mathcal{M}$ at $T$ enlarges the class of weak equivalences to incorporate the maps in $T$ while preserving the class of cofibrations. Consequently, the class of fibrations, and thus the class of fibrant objects, is forced to shrink. Though the fibrations in a left Bousfield localization can be quite difficult to describe, at least the fibrant objects are always known to us.
Theorem 7.3 ([2, Theorem 4.7]).
Let $\mathcal{M}$ be a combinatorial, left proper, and simplicial model category. Let $T$ be a set of morphisms in $\mathcal{M}$. There is a model structure $\mathcal{L}_{T}\mathcal{M}$ on the underlying category of $\mathcal{M}$, called the left Bousfield localization of $\mathcal{M}$ at $T$, in which a morphism $f$ is
@@
As the class of cofibrations is preserved, $\mathcal{L}_{T}\mathcal{M}$ inherits the generating cofibrations of the combinatorial model category $\mathcal{M}$. However, the same is not true for the generating acyclic cofibrations because the class of weak equivalences is expanded and, in general, it is difficult to describe the smaller class of fibrations in a left Bousfield localization. Knowing the fibrant objects in a left Bousfield localization makes it feasible to write down a set of maps that detects them by tweaking any set that detects fibrant objects in $\mathcal{M}$; however, this set can get cumbersome to write out, which defeats its practical purpose. In our left Bousfield localizations in this paper, we have no nice description of such a set that detects fibrant objects, so we avoid writing one out altogether.
@@ 7.2. The set for the left Bousfield localization
$\{S^{n}\rightarrow S^{n}_{M^{-1}\mathbb{Z}}\mid n\geq 2\}$
in order to left Bousfield localize the model category ${\mathcal{SS}\mathrm{ets}}_{1}$ at that set. A first suggestion is to start with the simplicial $n$-sphere $S_{\mathrm{simp}}^{n}$ and use it to emulate the construction of $S^{n}_{M^{-1}\mathbb{Z}}$ in the $1$-reduced setting. This approach runs into a problem: the construction of $S^{n}_{M^{-1}\mathbb{Z}}$ involves self-maps of $S^{n}$ of arbitrary degree, whereas, by the following lemma, the simplicial $n$-sphere $S_{\mathrm{simp}}^{n}$ only has the trivial self-maps.
@@
Let $f$ be a self-map of $S_{\mathrm{simp}}^{n}$. Let $x$ denote the 0-simplex of $S_{\mathrm{simp}}^{n}$, and let $\epsilon$ denote the non-degenerate $n$-simplex of $S_{\mathrm{simp}}^{n}$. If $f_{n}(\epsilon)=\epsilon$, then $f$ is the identity map of $S_{\mathrm{simp}}^{n}$. If $f_{n}(\epsilon)$ is the iterated degeneracy of $x$, then $f$ is the constant self-map of $S_{\mathrm{simp}}^{n}$. ∎
Thus, we have to settle for an indirect route. We start with the CW inclusion $S^{n}\rightarrow S^{n}_{M^{-1}\mathbb{Z}}$, and we apply the singular complex functor to obtain the inclusion of simplicial sets $\operatorname{Sing}(S^{n})\rightarrow\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}})$. Then, we take 11st Eilenberg subcomplexes to get the inclusion of $1$-reduced simplicial sets
@@
To confirm that this map is our desired analog, we study the commutative diagram of CW complexes
$|E_{1}\operatorname{Sing}(S^{n})|$$|\operatorname{Sing}(S^{n})|$$S^{n}$$|E_{1}\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}})|$$|\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}})|$$S^{n}_{M^{-1}\mathbb{Z}},$}}
where the right-hand horizontal maps arise from the geometric realization-singular complex adjunction.
@@
$|E_{1}\operatorname{Sing}(S^{n})|$$|\operatorname{Sing}(S^{n})|$$S^{n}$$|E_{1}\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}})|$$|\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}})|$$S^{n}_{M^{-1}\mathbb{Z}},$}}
@@
7.3. Description of the left Bousfield localization
@@
$B_{M^{-1}\mathbb{Z}}=\{E_{1}\operatorname{Sing}(S^{n})\rightarrow E_{1}\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}})\mid n\geq 2\}.$
In particular, we characterize the $B_{M^{-1}\mathbb{Z}}$-local objects and the $B_{M^{-1}\mathbb{Z}}$-local equivalences in ${\mathcal{SS}\mathrm{ets}}_{1}$. To state our characterizations, we produce analogs of the notions from Section 6.1 for $1$-reduced simplicial sets.
@@
    A $1$-reduced simplicial set $K$ is $M$-local if $|K|$ is an $M$-local space.
@@
    An $M$-localization of a $1$-reduced simplicial set $X$ is an $M$-local homotopy equivalence with $M$-local target $f\colon X\rightarrow X_{M^{-1}\mathbb{Z}}$.
@@
The following proposition helps us transition from $1$-reduced simplicial sets to $1$-connected spaces.
Proposition 7.8 ([26, Proposition 1.1.11]).
If $A$ and $X$ are pointed simplicial sets and $X$ is a Kan complex, then there is a natural weak equivalence $\operatorname{Map}_{\mathcal{SS}\mathrm{ets}_{\ast}}(A,X)\rightarrow\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}(|A|,|X|)$.
We show that a $1$-reduced simplicial set is $B_{M^{-1}\mathbb{Z}}$-local if and only if it is an $M$-local Kan complex.
@@
Let $M$ be a multiplicative subset of $\mathbb{Z}$. Let $X$ be a $1$-reduced Kan complex. Let $n\geq 2$. The following statements are equivalent.
@@
    The map $\operatorname{Map}_{\mathcal{SS}\mathrm{ets}_{\ast}}(E_{1}\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}}),X)\rightarrow\operatorname{Map}_{\mathcal{SS}\mathrm{ets}_{\ast}}(E_{1}\operatorname{Sing}(S^{n}),X)$ is a weak equivalence.
@@
    The map $\operatorname{Map}_{\mathcal{SS}\mathrm{ets}_{\ast}}(\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}}),X)\rightarrow\operatorname{Map}_{\mathcal{SS}\mathrm{ets}_{\ast}}(\operatorname{Sing}(S^{n}),X)$ is a weak equivalence.
@@
    The map $\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}(|\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}})|,|X|)\rightarrow\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}(|\operatorname{Sing}(S^{n})|,|X|)$ is a weak equivalence.
@@
    The map $\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}(S^{n}_{M^{-1}\mathbb{Z}},|X|)\rightarrow\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}(S^{n},|X|)$ is a weak equivalence.
@@
    The space $|X|$ is $M$-local.
@@
First, (1) and (2) are equivalent by Proposition 7.6. Then, (2) and (3) are equivalent by Proposition 7.8. Now, (3) and (4) are equivalent by Proposition 7.6. Lastly, (4) and (5) are equivalent by Proposition 6.9. ∎
By a similar series of characterizations, we establish that a map of $1$-reduced simplicial sets is a $B_{M^{-1}\mathbb{Z}}$-local equivalence if and only if it is an $M$-local homotopy equivalence. To that end, we need an auxiliary lemma.
@@
    For every $M$-local Kan complex $K$, the map $\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}(|f|,|K|)$ is a weak equivalence.
@@
    For every $M$-local pointed space $Z$, the map $\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}(|f|,Z)$ is a weak equivalence.
@@
Suppose (1) holds. Let $Z$ be an $M$-local pointed space. By Corollary 3.18, $E_{1}\operatorname{Sing}(Z)$ is a Kan complex. Corollary 4.4 and Theorem 4.1(2) provide the weak homotopy equivalences $|E_{1}\operatorname{Sing}(Z)|\rightarrow|\operatorname{Sing}(Z)|\rightarrow Z$, so $E_{1}\operatorname{Sing}(Z)$ is an $M$-local Kan complex. Now, there is a commutative diagram
$\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}(|Y|,|E_{1}\operatorname{Sing}(Z)|)$$\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}(|Y|,|\operatorname{Sing}(Z)|)$$\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}(|Y|,Z)$$\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}(|X|,|E_{1}\operatorname{Sing}(Z)|)$$\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}(|X|,|\operatorname{Sing}(Z)|)$$\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}(|X|,Z)$}}
whose horizontal maps are weak equivalences by Corollary 4.4 and Theorem 4.1(2) and whose left-hand vertical map is a weak equivalence by (1); thus, so is the right-hand vertical map, proving (2). Conversely, suppose (2) holds. Let $K$ be an $M$-local simplicial set. Then, $|K|$ is an $M$-local pointed space, which implies (1). ∎
Using our lemma, we show that a $B_{M^{-1}\mathbb{Z}}$-local equivalence is precisely an $M$-local homotopy equivalence.
@@
Let $M$ be a multiplicative subset of $\mathbb{Z}$. Let $f$ be a map of $1$-reduced simplicial sets. The following statements are equivalent.
@@
    For every $B_{M^{-1}\mathbb{Z}}$-local simplicial set $K$, the map $\operatorname{Map}_{\mathcal{SS}\mathrm{ets}_{\ast}}(f,K)$ is a weak equivalence.
@@
    For every $M$-local Kan complex $K$, the map $\operatorname{Map}_{\mathcal{SS}\mathrm{ets}_{\ast}}(f,K)$ is a weak equivalence.
@@
    For every $M$-local Kan complex $K$, the map $\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}(|f|,|K|)$ is a weak equivalence.